# {textblock}

### v1

@@ Abstract
Index Terms:
rdm, research data management, mde, model driven engineering, json, yaml, configuration, schema, data, model, editor, gui, tool, ai, llm, mapping, matching
@@ III-A AI-assisted JSON Schema Creation
For schema creation tasks, prompts are constructed according to the principles described above. Figure 1 shows an example of a user-provided schema description to MetaConfigurator and the resulting schema, visualized. The same chat-based interface can be used to modify existing schemas (Figure 2), to query a schema for information or to create, edit and query document instances.
@@ IV Application Example
Chemists have performed chemical synthesis experiments and tracked their inputs and results in an Excel table. They want to apply machine learning and other algorithms on their data, but it is not yet in a suitable format [^10] Sources: https://doi.org/10.18419/DARUS-5157.. MetaConfigurator is used to convert the excel table into a JSON document and automatically infer the corresponding schema. The JSON document and schema, however, are not yet of sufficient quality. Among other points, the product_purity values of yes and no have been inferred as being of type string instead of boolean and the ligand and metal_salt properties are all listed directly as children of an experiment instance, although it would be more suitable to extract them into separate ligand and metal_salt classes. Using the chat-based interface we create an improved schema (Figure 1-2). To transform the JSON document to satisfy the new schema, we use the automated schema mapping functionality. It successfully maps from yes and no to boolean values and from the flat source paths to the nested paths of the target schema. The data is now well-structured, formally specified by a schema and AI-ready.

### v2

# AI-assisted JSON Schema Creation and Mapping thanks: Deutsche Forschungsgemeinschaft (DFG) under project numbers 528693298 (preECO), 358283783 (SFB1333), and 390740016 (EXC2075)

@@ III-A AI-assisted JSON Schema Creation
For schema creation tasks, prompts are constructed according to the principles described above. Figure III-A shows an example of a user-provided schema description to MetaConfigurator and the resulting schema, visualized. The same chat-based interface can be used to modify existing schemas (Figure III-A), to query a schema for information or to create, edit and query document instances.
@@ IV Application Example
Chemists have performed chemical synthesis experiments and tracked their inputs and results in an Excel table. They want to apply machine learning and other algorithms on their data, but it is not yet in a suitable format [^10] Sources: https://doi.org/10.18419/DARUS-5157.. MetaConfigurator is used to convert the excel table into a JSON document and automatically infer the corresponding schema. The JSON document and schema, however, are not yet of sufficient quality. Among other points, the product_purity values of yes and no have been inferred as being of type string instead of boolean and the ligand and metal_salt properties are all listed directly as children of an experiment instance, although it would be more suitable to extract them into separate ligand and metal_salt classes. Using the chat-based interface we create an improved schema (Figure III-A-III-A). To transform the JSON document to satisfy the new schema, we use the automated schema mapping functionality. It successfully maps from yes and no to boolean values and from the flat source paths to the nested paths of the target schema. The data is now well-structured, formally specified by a schema and AI-ready.